# Scientists’ First Exam: Probing Cognitive Abilities of MLLM via Perception, Understanding, and Reasoning

### v5

@@ 1 Introduction
Scientific discoveries rely on investing significant time in analyzing large-scale, complex, and diverse data. Researchers require domain-specific knowledge to interpret scientific data across various modalities, and apply problem-solving skills to address specific scientific challenges [64]. Recent advances in multimodal large language models (MLLMs) have achieved remarkable performance on a wide range of benchmarks, comparable to or even surpassing human-level understanding in both general-level (e.g., MMLU [28], SuperGLUE [67], TriviaQA [31]) and graduate-level (e.g., GPQA [60], HumanEval [12], GSM8K [15]) knowledge domains. As MLLMs continue to progress from general-purpose understanding to domain-specific knowledge, scientific discovery has emerged as a critical frontier for evaluating and extending their abilities [5, 10, 11, 35, 42, 44, 84].
The process of scientific discovery often involves specialized scientific analysis of data modalities (e.g., molecular structures, spectra, protein sequences) from various scientific fields. SuperGPQA [17] extends conventional domains by incorporating long-tail disciplines, ensuring accessibility to real-world professional expertise. CURIE [16] establishes a ten-task benchmark for evaluating scientific reasoning in long-context scenarios. HLE [59] is introduced to evaluate model capabilities through challenging and expert-authored questions. However, despite the growing interest in the scientific domain, most existing scientific benchmarks extract tasks from secondary sources such as academic materials [16, 27, 78] and textbooks [70, 78]. As a result, they inadequately probe the cognitive abilities (e.g., perception, understanding, and reasoning) required for analyzing scientific data encountered in real-world research. Moreover, these benchmarks tend to focus exclusively on a single ability to interpret domain knowledge from the data, while neglecting the full spectrum from perception to reasoning. This capability gap exposes a fundamental challenge: How to granularly measure MLLMs’ scientific cognitive capabilities across multiple disciplines for scientific discovery?
@@
We benchmark 16 state-of-the-art open and closed weight MLLMs using SFE, as illustrated in Fig. 1(c). As observed, while these MLLMs perform well on benchmarks such as MMLU [78] and ScienceQA [48], they all exhibit suboptimal results on the SFE benchmark. This indicates that SFE serves as a challenging frontier for scientific MLLM development. Our contributions:
@@ 2 Related Works
MMMU[78] AD, Bus., Sci. Med., HSS, TE MCQ OQ .png Textbooks e-Resources EN
ScienceQA[48] Natural, social and language science MCQ .png Curricula EN
SciBench[70] Math, Chem Physics OQ .png Textbooks EN
CMMU[27] 7 School subjects OQ .png Exams ZH
ChemBench[26] Chem EM - 8 Public datasets EN
SuperGPQA[17] 13 Disciplines MCQ - Experts EN
SciEval[63] Chem, Phy, Bio MCQ, EM judgment - Knowledge base EN
HLE[59] Math, Phy, Bio HSS, CS, Chem,TE MCQ, EM .png Experts EN
CURIE[16] Mat, Phy, QC, Geo Bio, Proteins OQ .png 10 Experts EN
@@
Science LLMs / MLLMs. Recent advancements in domain-specific large language models (LLMs) have significantly impacted various scientific fields. In the biomedical domain, LLMs have been employed for tasks such as clinical documentation, information retrieval, and hypothesis generation  [35, 36, 49, 50, 51, 56]. In chemistry, recent studies have focused on tasks such as drug property prediction, molecular discovery, chemical reaction extraction, and protein structure understanding [11, 39, 40, 43, 44, 45, 47, 52, 81, 87]. Notably, researchers have developed LLMs trained on scientific corpora to support molecular and protein discovery  [5, 14, 20, 54, 66, 71]. For instance, iupacGPT [14] uses IUPAC nomenclature to effectively capture relationships among atoms and chemical groups. Similarly, Progen  [54] and ProtGPT2  [20] are trained for protein sequence generation and understanding. Geosciences have benefited from LLMs through their applications in ocean science, extreme weather, and remote sensing [8, 38, 41, 53, 73, 77, 83, 84]. Models like EarthGPT [83] and CLLMate [38] further integrate multimodal knowledge to support scientific question answering.
Science Benchmarks. With the development of large language models (LLMs), recent efforts in LLM benchmarking have increasingly focused on evaluating scientific reasoning capabilities across diverse domains and modalities. Early benchmarks [46] such as ScienceQA [48] and CMMU [27], primarily focus on science level below high school. Several college-level benchmarks [17, 70, 72, 78, 79] have emerged to support deeper scientific understanding. As state-of-the-art multimodal LLMs rapidly improve, many of them now achieve strong performance on these benchmarks. This has motivated the creation of more expert-level datasets that feature high-difficulty questions such as HLE [59] and CURIE [16]. However, although some benchmarks offer broad domain coverage, they often lack clearly defined tasks, making it difficult to assess specific model limitations. To address this, works like ChemLLMBench [26] and CURIE [16] have introduced smaller-scale but task-oriented benchmarks that enable more targeted evaluation of scientific capabilities. Nevertheless, nearly all of these benchmarks remain monolingual, limiting their applicability for comprehensively evaluating LLMs in global deployment contexts.
@@ 3.1 Data Collection
Task Construction. Building upon the established scientific directions, domain experts collaboratively define concrete tasks. We first formalize the task definition rules by mapping each scientific direction to an appropriate cognitive level and question type as shown in Fig. 2. This taxonomy ensures that each task is aligned with a distinct level of cognitive capacity and supports evaluation diversity. For each discipline, we collaborated with 2-5 experts to define benchmark tasks that align with high-priority issues in real-world scientific research. Each expert-designed task is expected to satisfy the following core criteria: (1) It reflects a meaningful problem that requires substantial domain knowledge and is commonly encountered in real-world research. (2) It must be solvable through expert-level reasoning that necessarily integrates multimodal scientific inputs such as structured visualizations and symbolic representations. (3) It aligns with one of the proposed cognitive levels (L1-L3). To support each task, experts also provide instructions for collecting raw data that would serve as visual inputs for the VQA pairs. These data sources span a diverse range of formats, including public datasets (e.g., ERA5 [29]), existing benchmarks (e.g., MoleculeNet [75]), and domain-specific databases (e.g., RCSB PDB [7], PDBbind [69], and PubChem [33]). For each task, domain experts identify suitable data entries that are typically used to analyze the corresponding scientific problems, such as RNA sequence files (.stockholm) and protein structure files (.pdb).
@@ 4 Experiments and Evaluations
General Settings. We conduct a comprehensive evaluation of the state-of-the-art MLLMs on SFE. For models with open weights, we assess InternVL2.5-78B [13], InternVL3-78B [86], Qwen2.5-VL-72B [6], LLaMaVision-90B [22], and LLaVa-Onevision-72B [37]. For models with closed weights, we evaluate GPT-4o-2024-11-20 [57], GPT-4.1-2025-04-14 [57], GPT-o1-2024-12-17 [57], GPT-o3 [57], Claude 3 Opus [4], Claude 3.7 Sonnet [4], Gemini-2.0-Flash [65], Gemini-2.5-Flash [65], Gemini-2.5-Pro [65], Grok-2-Vision-12-12 [76] and Doubao-1.5-Vision-Pro [24]. When benchmarking, we configure all MLLMs’ temperatures to 0 for reduced randomness and employ a standard zero-shot prompt template across all tasks. Specifically, the template begins with a description of the task assigned to the model, followed by the inclusion of question texts with interleaved images. Additionally, we fix the maximal number of generated tokens to 1024, ensuring fairness and cost-effectiveness in our evaluations.
Metrics. We present the BERTScore [82] and the LLM-as-a-Judge score [23] for all tasks, except for the remote sensing perception task in Earth science, where we report the execution success rate and the Intersection over Union (IoU). For BERTScore, we use the F1 score. For LLM-as-a-Judge score, we employ GPT-4o-2024-11-20 as the judge, allowing us to semantically verify the correctness of answers against model predictions. Conversely, the execution success rate is utilized to assess whether MLLMs can accurately follow prompts to produce bounding boxes in the desired format. Finally, the IoU metric evaluates the precision of these bounding boxes compared to ground truths. Without further clarification, each experiment is conducted once to obtain the final results.
@@ 4.2 Analysis
Pass@k Analysis. The Pass@k metric [34] selects the highest quality answers from an MLLM as its final response to a question, indicating the model’s potential for improvement through post-training (e.g., RLHF [58], GRPO [25], etc.). In Fig. 3, we evaluate pass@k scores of GPT-4.1-2025-04-14, Gemini-2.5-Flash, Qwen2.5-VL-72b and InternVL-3-78B on SFE, with $k$ ranging from $1$ to $6$. As shown, GPT-4.1-2025-04-14 and Gemini-2.5-Flash outperform state-of-the-art open-weight MLLMs by not only exhibiting superior initial performance (30.56% v.s. 26.09%) but also demonstrating strong scalability (30.56% $\rightarrow$ 37.75% v.s. 26.09 $\rightarrow$ 27.33%). This suggests that closed-weight MLLMs may leverage more diverse and expansive raw datasets during pre-training than open-weight MLLMs. Furthermore, their post-training phase may prioritize a balanced approach, emphasizing exploration alongside exploitation, rather than focusing exclusively on exploitation [80].
@@ Appendix C Tasks Description
Astronomy. In the field of astronomy, raw astronomical observations span multiple modalities, including images, time series, spectra, and hyperspectral observations [1, 32, 2]. Specifically, data collected for celestial objects often covers multiple bands and includes rich attribute information. Therefore, inspired by [3], we divide astronomical tasks into three categories based on data modalities, including Celestial Image Understanding and Comparison, Time-series Light Curve Analysis, and Spectral Analysis, each of which is detailed below.
@@
  • Structure Analysis. Structure Analysis tests a model’s grasp of symbolic chemistry, such as identifying elemental composition, describing molecules using natural language, and generating IUPAC (International Union of Pure and Applied Chemistry) names (e.g., methane, ethane) [55]. It examines the foundational ability to ensure that the model correctly recognizes the chemical structure and elements of the molecules.
  • Property Prediction. Molecular Property Prediction aims to assess a model’s ability to infer chemical and biological properties from molecular structures. It focuses on the properties of chemical molecules that scientists are most interested in, such as drug molecules’ ADMET (i.e., absorption, distribution, metabolism, excretion, toxicity) [19], synthetic accessibility, and quantum mechanics energy, etc.
@@
  • Molecular Design and Optimization. Molecular Design and Optimization evaluates the MLLM’s capacity to generate molecules with desirable chemical properties (e.g., pharmaceutical properties). These tasks focus on identifying novel and diverse molecules with desirable molecular properties (e.g., druglikeness, solubility, binding affinities to the target proteins), which are fundamental tasks in drug and material discovery [18].
  • Drug Screening. Drug Screening evaluates the MLLM’s ability to identify biologically relevant candidate molecules based on protein sequences and molecular structure information [30].
Earth. The Earth science domain includes various data modalities. First, variables can be either temporal or static. Meteorological time series are often used to analyze strong convective weather events  [85, 21, 74]. Static attributes, such as long-term average temperature maps, are typically used to capture climate patterns. Second, strong correlations often exist between variables. For example, the position of the western Pacific subtropical high (i.e., the 5880 hPa line) is positively correlated with the coverage range of precipitation belts. Third, variables may originate from diverse sources, including ground-based weather radar  [62], multi-spectral satellite imagers  [61], and various remote sensing instruments. To cater to such diverse scenarios, we establish a comprehensive Earth science task paradigm that focuses on 4 tasks, each comprising multiple sub-tasks.
@@
  • Task4: Ground Data Comparison Reasoning. Ground observational data have been widely used in scenarios such as global climatological analysis and extreme precipitation monitoring  [68, 9]. To evaluate the capability of MLLMs in comparing and reasoning meteorological terminology and data distributions, we define 4 tasks based on different variables. For example, in the temperature sequence comparison task, MLLMs need to compare annual temperature series from two different time periods and describe differences of statistical characteristics

### v6

@@ 1 Introduction
Scientific discoveries rely on investing significant time in analyzing large-scale, complex, and diverse data. Researchers require domain-specific knowledge to interpret scientific data across various modalities, and apply problem-solving skills to address specific scientific challenges [taconis2001teaching]. Recent advances in multimodal large language models (MLLMs) have achieved remarkable performance on a wide range of benchmarks, comparable to or even surpassing human-level understanding in both general-level (e.g., MMLU [hendrycks2020measuring], SuperGLUE [wang2019superglue], TriviaQA [joshi2017triviaqa]) and graduate-level (e.g., GPQA [rein2024gpqa], HumanEval [chen2021evaluating], GSM8K [cobbe2021training]) knowledge domains. As MLLMs continue to progress from general-purpose understanding to domain-specific knowledge, scientific discovery has emerged as a critical frontier for evaluating and extending their abilities [bagal2021molgpt, cao2023instructmol, chen2024pharmagpt, labrak2024biomistral, liu2024git, liu2024drugllm, zhang2025earthgpt].
The process of scientific discovery often involves specialized scientific analysis of data modalities (e.g., molecular structures, spectra, protein sequences) from various scientific fields. SuperGPQA [du2025supergpqa] extends conventional domains by incorporating long-tail disciplines, ensuring accessibility to real-world professional expertise. CURIE [cuicurie] establishes a ten-task benchmark for evaluating scientific reasoning in long-context scenarios. HLE [phan2025humanity] is introduced to evaluate model capabilities through challenging and expert-authored questions. However, despite the growing interest in the scientific domain, most existing scientific benchmarks extract tasks from secondary sources such as academic materials [cuicurie, he2024cmmu, yue2024mmmu] and textbooks [wang2024scibench, yue2024mmmu]. As a result, they inadequately probe the cognitive abilities (e.g., perception, understanding, and reasoning) required for analyzing scientific data encountered in real-world research. Moreover, these benchmarks tend to focus exclusively on a single ability to interpret domain knowledge from the data, while neglecting the full spectrum from perception to reasoning. This capability gap exposes a fundamental challenge: How to granularly measure MLLMs’ scientific cognitive capabilities across multiple disciplines for scientific discovery?
@@
We benchmark 16 state-of-the-art open and closed weight MLLMs using SFE, as illustrated in Fig. 1(c). As observed, while these MLLMs perform well on benchmarks such as MMLU [yue2024mmmu] and ScienceQA [lu2022learn], they all exhibit suboptimal results on the SFE benchmark. This indicates that SFE serves as a challenging frontier for scientific MLLM development. Our contributions:
@@ 2 Related Works
MMMU[yue2024mmmu] AD, Bus., Sci. Med., HSS, TE MCQ OQ .png Textbooks e-Resources EN
ScienceQA[lu2022learn] Natural, social and language science MCQ .png Curricula EN
SciBench[wang2024scibench] Math, Chem Physics OQ .png Textbooks EN
CMMU[he2024cmmu] 7 School subjects OQ .png Exams ZH
ChemBench[guo2023can] Chem EM - 8 Public datasets EN
SuperGPQA[du2025supergpqa] 13 Disciplines MCQ - Experts EN
SciEval[sun2024scieval] Chem, Phy, Bio MCQ, EM judgment - Knowledge base EN
HLE[phan2025humanity] Math, Phy, Bio HSS, CS, Chem,TE MCQ, EM .png Experts EN
CURIE[cuicurie] Mat, Phy, QC, Geo Bio, Proteins OQ .png 10 Experts EN
@@
Science LLMs / MLLMs. Recent advancements in domain-specific large language models (LLMs) have significantly impacted various scientific fields. In the biomedical domain, LLMs have been employed for tasks such as clinical documentation, information retrieval, and hypothesis generation  [labrak2024biomistral, lee2020biobert, luo2022biogpt, luo2023biomedgpt, luu2024bioinspiredllm, miolo2021electramed]. In chemistry, recent studies have focused on tasks such as drug property prediction, molecular discovery, chemical reaction extraction, and protein structure understanding [chen2024pharmagpt, li2024empowering, liang2023drugchat, liu2024conversational, liu2024drugllm, liu2024moleculargpt, liu2023molxpt, lv2025prollama, zhang2024chemllm, zhuo2024protllm]. Notably, researchers have developed LLMs trained on scientific corpora to support molecular and protein discovery  [bagal2021molgpt, cho2023iupacgpt, ferruz2022protgpt2, madani2020progen, truong2023poet, wang2023cmolgpt]. For instance, iupacGPT [cho2023iupacgpt] uses IUPAC nomenclature to effectively capture relationships among atoms and chemical groups. Similarly, Progen  [madani2020progen] and ProtGPT2  [ferruz2022protgpt2] are trained for protein sequence generation and understanding. Geosciences have benefited from LLMs through their applications in ocean science, extreme weather, and remote sensing [bi2023oceangpt, li2024cllmate, lin2023geogalactica, ma2024weatherqa, webersinke2021climatebert, yan2024urbanclip, zhang2024earthgpt, zhang2025earthgpt]. Models like EarthGPT [zhang2024earthgpt] and CLLMate [li2024cllmate] further integrate multimodal knowledge to support scientific question answering.
Science Benchmarks. With the development of large language models (LLMs), recent efforts in LLM benchmarking have increasingly focused on evaluating scientific reasoning capabilities across diverse domains and modalities. Early benchmarks [liu2024mmbench] such as ScienceQA [lu2022learn] and CMMU [he2024cmmu], primarily focus on science level below high school. Several college-level benchmarks [du2025supergpqa, wang2024scibench, wang2024mmlu, yue2024mmmu, yue2024mmmu2] have emerged to support deeper scientific understanding. As state-of-the-art multimodal LLMs rapidly improve, many of them now achieve strong performance on these benchmarks. This has motivated the creation of more expert-level datasets that feature high-difficulty questions such as HLE [phan2025humanity] and CURIE [cuicurie]. However, although some benchmarks offer broad domain coverage, they often lack clearly defined tasks, making it difficult to assess specific model limitations. To address this, works like ChemLLMBench [guo2023can] and CURIE [cuicurie] have introduced smaller-scale but task-oriented benchmarks that enable more targeted evaluation of scientific capabilities. Nevertheless, nearly all of these benchmarks remain monolingual, limiting their applicability for comprehensively evaluating LLMs in global deployment contexts.
@@ 3.1 Data Collection
Task Construction. Building upon the established scientific directions, domain experts collaboratively define concrete tasks. We first formalize the task definition rules by mapping each scientific direction to an appropriate cognitive level and question type as shown in Fig. 2. This taxonomy ensures that each task is aligned with a distinct level of cognitive capacity and supports evaluation diversity. For each discipline, we collaborated with 2-5 experts to define benchmark tasks that align with high-priority issues in real-world scientific research. Each expert-designed task is expected to satisfy the following core criteria: (1) It reflects a meaningful problem that requires substantial domain knowledge and is commonly encountered in real-world research. (2) It must be solvable through expert-level reasoning that necessarily integrates multimodal scientific inputs such as structured visualizations and symbolic representations. (3) It aligns with one of the proposed cognitive levels (L1-L3). To support each task, experts also provide instructions for collecting raw data that would serve as visual inputs for the VQA pairs. These data sources span a diverse range of formats, including public datasets (e.g., ERA5 [hersbach2020era5]), existing benchmarks (e.g., MoleculeNet [wu2018moleculenet]), and domain-specific databases (e.g., RCSB PDB [berman2003announcing], PDBbind [wang2004pdbbind], and PubChem [kim2024pubchem]). For each task, domain experts identify suitable data entries that are typically used to analyze the corresponding scientific problems, such as RNA sequence files (.stockholm) and protein structure files (.pdb).
@@ 4 Experiments and Evaluations
General Settings. We conduct a comprehensive evaluation of the state-of-the-art MLLMs on SFE. For models with open weights, we assess InternVL2.5-78B [chen2024expanding], InternVL3-78B [zhu2025internvl3], Qwen2.5-VL-72B [Qwen2.5-VL], LLaMaVision-90B [grattafiori2024llama], and LLaVa-Onevision-72B [li2024llava]. For models with closed weights, we evaluate GPT-4o-2024-11-20 [gpt2023], GPT-4.1-2025-04-14 [gpt2023], GPT-o1-2024-12-17 [gpt2023], GPT-o3 [gpt2023], Claude 3 Opus [anthropic2023], Claude 3.7 Sonnet [anthropic2023], Gemini-2.0-Flash [team2023gemini], Gemini-2.5-Flash [team2023gemini], Gemini-2.5-Pro [team2023gemini], Grok-2-Vision-12-12 [grok2023] and Doubao-1.5-Vision-Pro [guo2025seed1]. When benchmarking, we configure all MLLMs’ temperatures to 0 for reduced randomness and employ a standard zero-shot prompt template across all tasks. Specifically, the template begins with a description of the task assigned to the model, followed by the inclusion of question texts with interleaved images. Additionally, we fix the maximal number of generated tokens to 1024, ensuring fairness and cost-effectiveness in our evaluations.
Metrics. We present the BERTScore [zhang2019bertscore] and the LLM-as-a-Judge score [gu2024survey] for all tasks, except for the remote sensing perception task in Earth science, where we report the execution success rate and the Intersection over Union (IoU). For BERTScore, we use the F1 score. For LLM-as-a-Judge score, we employ GPT-4o-2024-11-20 as the judge, allowing us to semantically verify the correctness of answers against model predictions. Conversely, the execution success rate is utilized to assess whether MLLMs can accurately follow prompts to produce bounding boxes in the desired format. Finally, the IoU metric evaluates the precision of these bounding boxes compared to ground truths. Without further clarification, each experiment is conducted once to obtain the final results.
@@ 4.2 Analysis
Pass@k Analysis. The Pass@k metric [kulal2019spoc] selects the highest quality answers from an MLLM as its final response to a question, indicating the model’s potential for improvement through post-training (e.g., RLHF [ouyang2022training], GRPO [guo2025deepseek], etc.). In Fig. 3, we evaluate pass@k scores of GPT-4.1-2025-04-14, Gemini-2.5-Flash, Qwen2.5-VL-72b and InternVL-3-78B on SFE, with $k$ ranging from $1$ to $6$. As shown, GPT-4.1-2025-04-14 and Gemini-2.5-Flash outperform state-of-the-art open-weight MLLMs by not only exhibiting superior initial performance (30.56% v.s. 26.09%) but also demonstrating strong scalability (30.56% $\rightarrow$ 37.75% v.s. 26.09 $\rightarrow$ 27.33%). This suggests that closed-weight MLLMs may leverage more diverse and expansive raw datasets during pre-training than open-weight MLLMs. Furthermore, their post-training phase may prioritize a balanced approach, emphasizing exploration alongside exploitation, rather than focusing exclusively on exploitation [yue2025does].
@@ Appendix C Tasks Description
Astronomy. In the field of astronomy, raw astronomical observations span multiple modalities, including images, time series, spectra, and hyperspectral observations [aghamousa2016desi, kessler2019models, aihara2018hyper]. Specifically, data collected for celestial objects often covers multiple bands and includes rich attribute information. Therefore, inspired by [angeloudi2024multimodal], we divide astronomical tasks into three categories based on data modalities, including Celestial Image Understanding and Comparison, Time-series Light Curve Analysis, and Spectral Analysis, each of which is detailed below.
@@
  • Structure Analysis. Structure Analysis tests a model’s grasp of symbolic chemistry, such as identifying elemental composition, describing molecules using natural language, and generating IUPAC (International Union of Pure and Applied Chemistry) names (e.g., methane, ethane) [mcnaught2006iupac]. It examines the foundational ability to ensure that the model correctly recognizes the chemical structure and elements of the molecules.
  • Property Prediction. Molecular Property Prediction aims to assess a model’s ability to infer chemical and biological properties from molecular structures. It focuses on the properties of chemical molecules that scientists are most interested in, such as drug molecules’ ADMET (i.e., absorption, distribution, metabolism, excretion, toxicity) [ferreira2019admet], synthetic accessibility, and quantum mechanics energy, etc.
@@
  • Molecular Design and Optimization. Molecular Design and Optimization evaluates the MLLM’s capacity to generate molecules with desirable chemical properties (e.g., pharmaceutical properties). These tasks focus on identifying novel and diverse molecules with desirable molecular properties (e.g., druglikeness, solubility, binding affinities to the target proteins), which are fundamental tasks in drug and material discovery [du2024machine].
  • Drug Screening. Drug Screening evaluates the MLLM’s ability to identify biologically relevant candidate molecules based on protein sequences and molecular structure information [huang2022artificial].
Earth. The Earth science domain includes various data modalities. First, variables can be either temporal or static. Meteorological time series are often used to analyze strong convective weather events  [zhang2023skilful, gong2024cascast, wilson1998nowcasting]. Static attributes, such as long-term average temperature maps, are typically used to capture climate patterns. Second, strong correlations often exist between variables. For example, the position of the western Pacific subtropical high (i.e., the 5880 hPa line) is positively correlated with the coverage range of precipitation belts. Third, variables may originate from diverse sources, including ground-based weather radar  [smith2016multi], multi-spectral satellite imagers  [schmit2010goes], and various remote sensing instruments. To cater to such diverse scenarios, we establish a comprehensive Earth science task paradigm that focuses on 4 tasks, each comprising multiple sub-tasks.
@@
  • Task4: Ground Data Comparison Reasoning. Ground observational data have been widely used in scenarios such as global climatological analysis and extreme precipitation monitoring  [wang2021characteristic, binetti2022use]. To evaluate the capability of MLLMs in comparing and reasoning meteorological terminology and data distributions, we define 4 tasks based on different variables. For example, in the temperature sequence comparison task, MLLMs need to compare annual temperature series from two different time periods and describe differences of statistical characteristics